OpenStudy (anonymous):

okay so, how can i verify that cos(x)/(1-sin(x))=sec(x)+tan(x)? I changed the right side to other terms, but then i just came out with cos(x)^2 on the left and zero on the right. any ideas?

OpenStudy (anonymous):

square everything and then use the property rules

jimthompson5910 (jim_thompson5910):

cos(x)/(1-sin(x))=sec(x)+tan(x) [cos(x)/(1-sin(x))][(1+sin(x))/(1+sin(x))] = sec(x)+tan(x) [cos(x)(1+sin(x))]/[(1-sin(x))(1+sin(x))] = sec(x)+tan(x) (cos(x)+sin(x)cos(x))/(1-sin^2(x)) = sec(x)+tan(x) (cos(x)+sin(x)cos(x))/(cos^2(x)) = sec(x)+tan(x) (cos(x)/(cos^2(x))+(sin(x)cos(x))/(cos^2(x)) = sec(x)+tan(x) 1/cos(x)+sin(x)/cos(x) = sec(x)+tan(x) sec(x)+tan(x) = sec(x)+tan(x)